### v1

# HGRN2: Gated Linear RNNs with State Expansion

Abstract
Hierarchically gated linear RNN (HGRN, Qin et al. 2023c) has demonstrated competitive training speed and performance in language modeling, while offering efficient inference. However, the recurrent state size of HGRN remains relatively small, which limits its expressiveness. To address this issue, inspired by linear attention, we introduce a simple outer-product-based state expansion mechanism so that the recurrent state size could be significantly enlarged without introducing any additional parameters. The linear attention form also allows for hardware-efficient training. Our extensive experiments verify the advantage of HGRN2 over HGRN1 in language modeling, image classification, and Long Range Arena. Our largest 3B HGRN2 model slightly outperforms Mamba and LLaMa Architecture Transformer for language modeling in a controlled experiment setting; and performs competitively with many open-source 3B models in downstream evaluation but using much fewer total training tokens.
1 Introduction
Large language models (LLMs) have achieved great empirical success in recent years. However, serving Transformer-based LLMs is expensive due to the expensive KV cache management. Recurrent neural networks (RNNs), on the other hand, offer linear inference complexity with constant state size, which is ideal for serving. Therefore, there is great interest in studying parallelizable linear recurrent models such as linear RNNs (Peng et al., 2023; Orvieto et al., 2023; Qin et al., 2023c; De et al., 2024), linear attention (Sun et al., 2023b; Qin et al., 2023b; Yang et al., 2023; Arora et al., 2024; Peng et al., 2024), and state space models (Gu et al., 2022a; Smith et al., 2023; Gu & Dao, 2023).
RNNs have a fixed recurrent state size to encode all historical information. Therefore, it is important for RNNs (i) to better exploit the fix-sized states and (ii) to increase the recurrent state size. Recent improvements in linear RNN are exactly on this track, including the use of data-dependent decays and state expansion techniques.
Data-dependent decays (also known as forget gates) are of great importance for RNNs (van der Westhuizen & Lasenby, 2018), which allows them to selectively retain useful information while erasing irrelevant information, so that the fixed-size recurrent state could be better exploited. HGRN (Qin et al., 2023c) first highlighted the importance of data-dependent decays for linear RNNs. Since then, many recent linear recurrent models have also employed data-dependent decays, such as Mamba (Gu & Dao, 2023), Gated Linear Attention (GLA, Yang et al. 2023), Griffin (De et al., 2024), and RWKV-6 (Peng et al., 2024).
However, HGRN did not increase the recurrent state size size, greatly restricted by the limited memory capacity, which prevents it from achieving LLaMa-like (Touvron et al., 2023a; b) language modeling performance, as noted in Qin et al. (2024). Recent state-of-the-art linear recurrent models, such as Mamba, GLA, and RWKV-6, additionally use state expansion techniques, which greatly increase the recurrent state size and thereby enhance the memory capacity.
In this work, we propose HGRN2, with the aim of increasing the size of the recurrent state for HGRN while retaining both parameter efficiency and training efficiency. We first explore structured matrices to directly expand the state size in a parameter-efficient manner. Empirically, we found that they improved the language modeling performance, but still faced with the training inefficiency issue, thus they cannot scale the recurrent state size well. Inspired by linear attention, we then explore using a nonparameteric outer product-based state expansion mechanism, which allows efficient scaling of the size of the recurrent state during training without introducing any additional parameters. Thanks to the gated linear attention form, we could borrow the hardware-efficient GLA training algorithm described in Yang et al. (2023) for large-scale experiments.
We extensively evaluate HGRN2 in language modeling, image classification, and the Long Range Arena benchmark. HGRN2 outperforms HGRN1 in all these benchmarks, showing the significant benefit of using state expansion. Our largest 3B HGRN2 model slightly outperforms Mamba and LLaMa Architecture Transformer for language modeling in a controlled experiment setting; and performs competitively with many open-source 3B models in downstream evaluation but using much fewer total training tokens.
2 Background
2.1 Gated linear RNN
Given input $\mathbf{x}\in\mathbb{R}^{N\times d}$ where the sequence length is $N$ and the model dimension is $d$. A minimalist gated linear recurrent layer transforms the input $\mathbf{x}$ into hidden states $\mathbf{h}\in\mathbb{R}^{N\times d}$ and the output $\mathbf{y}\in\mathbb{R}^{N\times d}$, as defined below.
$\mathbf{g}_{t}$ $=\sigma\left(\mathbf{U}\mathbf{x}_{t}+\mathbf{b}_{u}\right),$ (1)
$\mathbf{i}_{t}$ $=\mathbf{V}\mathbf{x}_{t}+\mathbf{b}_{v},$
$\mathbf{o}_{t}$ $=\tau\left(\mathbf{W}\mathbf{x}_{t}+\mathbf{b}_{w}\right),$
$\mathbf{h}_{t}$ $=\mathbf{g}_{t}\odot\mathbf{h}_{t-1}+\left(1-\mathbf{g}_{t}\right)\odot\mathbf{i}_{t},$
$\mathbf{y}_{t}$ $=\mathbf{h}_{t}\odot\mathbf{o}_{t},$
where $\odot$ is element-wise product; $\sigma$ is the sigmoid function and $\tau$ is nonlinear activation (we choose to use $\mathrm{SiLU}$); $\mathbf{i}_{t}$ is the input vector; $\mathbf{g}_{t},\mathbf{o}_{t}$ are the forget gate and output gate, respectively. The input gate is tied to the forget gate: $1-\mathbf{g}_{t}$, which is commonly used in many gated RNNs such as GRU (Chung et al., 2014).
2.2 HGRN (Qin et al., 2023c)
Compared to Eq. 1, HGRN makes two adjustments: (i) complex-valued recurrence. (ii) forget gates with monotonically increased lower bound values from bottom layers to upper layers.
For (i), similarly to the findings of Gu & Dao (2023) and De et al. (2024), we empirically found that the complex-valued recurrence is not necessary, as shown in Table 1. We speculate that the reason Qin et al. (2023c) found it useful is due to state expansion: The complex-valued recurrent state is twice the size of that in the real-valued recurrent state. If we directly expand the real-valued recurrent state size from $d$ to $2d$, the language modeling performance on Wikitext-103 corpus is even better. Therefore, we only consider the real-valued recurrence thereafter.
Table 1: Comparison of real HGRN and complex HGRN. We found that real HGRN with twice the state size performs better than complex HGRN in Wiki103 language modeling.
Method State size PPL(val) PPL(test) Params (M)
Complex HGRN1 $2d$ 24.14 24.82 46.25
Real HGRN1 $d$ 25.34 26.12 46.24
Real HGRN1 $2d$ 23.81 24.55 47.82
For (ii), suppose that the total number of layers is $L$, and HGRN introduces a data-independent learnable matrix $\Gamma\in\mathbb{R}^{L\times d}$. $\Gamma_{i}$ will be the lowest values of the forget gate for the $i$th layer in all time steps. HGRN argues that this lower bound should be monotonically increasing from bottom to top, encouraging the bottom layers to model short-term local dependencies and the upper layers to model long-term dependencies. To enforce such monotonicity, HGRN used the cumulative softmax operator cumax [^1] This was first introduced by Shen et al. (2018):
$\beta:=\texttt{cumax}(\Gamma)=\texttt{cumsum}(\texttt{softmax}(\Gamma,\mathrm{dim=0}),\mathrm{dim=0})\in\mathbb{R}^{L\times d},\beta^{i}=[\beta]_{i}\in\mathbb{R}^{d}.$
To avoid the lower bound being one in the highest layer, HGRN subtracts all $\beta$ by $\beta^{0}$, so that the lower bound for the first layer is zero. After obtaining the lower bound values, the old forget gate $\mathbf{g}_{t}$ learns the residuals instead, resulting the new forget gate $\mathbf{f}_{t}$
$\mathbf{f}_{t}^{i}$ $={\beta}^{i}+(1-{\beta}^{i})\odot\mathbf{g}_{t}^{i},$ (2)
$\mathbf{h}_{t}^{i}$ $=\mathbf{f}_{t}^{i}\odot\mathbf{h}_{t-1}^{i}+(1-\mathbf{f}_{t}^{i})\odot\mathbf{i}_{t}^{i},$
where we use superscript to indicate the layer index.
Figure 1: The neural architecture of HGRN2. Each HGRN2 layer includes a token mixer layer HGRU2 and a channel mixerlayer GLU. HGRU2 employs recurrent computation through Eq. 3, where $\mathbf{i}_{t}$ is the input vector, $\mathbf{g}_{t}$ is the forget gate (not lower bounded), $\beta^{i}$ is the lower bound of the forget gate value, $\mathbf{o}_{t}$ is the output gate for layer $i$.
3 Method
3.1 Explorations of state expansion methods
The goal is to scale the size of the HGRN recurrent state from $d$ to $nd$, where $n$ is the expansion ratio of the state. However, if we use the original parameterization in Eq. 1, $\mathbf{U},\mathbf{V},\mathbf{W}$ will be of size $d\times nd$, which is very parameter inefficient when $n$ is large. Ideally, the number of parameters should be around $d^{2}$ as in the original case for each projection. First, we consider the following structured matrices, including low-rank, group linear transformation (GLT, Mehta et al. 2018; 2021), group linear transformation with group interaction (GLTI), Khatri-Rao product (KRP) and Kronecker product (KP), to replace the dense projection matrix $\mathbb{R}^{d}\rightarrow\mathbb{R}^{nd}$ as described in Table 2.
Table 2: We use Einstein Summation notation to represent Parameter Efficient State Expand (PESE) methods. We use Blue to represent the input, Black to represent data independent weight and Red to represent the output.
PESE Naive LR GLT
Equation ${{\color[rgb]{0,0,1}d},d\ nd\to{\color[rgb]{1,0,0}nd}}$ ${\color[rgb]{0,0,1}d},d\ r,r\ nd\to{\color[rgb]{1,0,0}nd}$ $\color[rgb]{0,0,1}d=(n\ e)\to n\ e$ ${\color[rgb]{0,0,1}n\ e},n\ e\ d\to{\color[rgb]{1,0,0}(n\ d)}$
Prameters $nd^{2}$ $dr(n+1)\approx d^{2}$ $d^{2}$
PESE GLTI KRP KP
Equation $\color[rgb]{0,0,1}d=(n\ e)\to n\ e$ ${\color[rgb]{0,0,1}n^{\prime}\ e},n^{\prime}\ e\ d,n^{\prime}\ n\to{\color[rgb]{1,0,0}(n\ d)}$ ${\color[rgb]{0,0,1}d},n\ d\to{\color[rgb]{1,0,0}(n\ d)}$ ${\color[rgb]{0,0,1}d},d\ d,n\to{\color[rgb]{1,0,0}(n\ d)}$
Parameters $d^{2}+n^{2}$ $nd$ $d^{2}+n$
After obtaining expanded $\mathbf{g},\mathbf{i},\mathbf{o}$, we feed them into element-wise gated linear recurrent layers as in Eq 1 and  2, resulting in the output vector $\mathbf{y}_{t}\in\mathbb{R}^{n\times d}$. To project the expanded dimension back to the original dimension, we simply marginalize the dimension corresponding to $n$.
The result is shown in Table 3. We found that state expansion generally improves the performance and that the low-rank matrix performs the best among these candidates.
However, these PESE methods face the training inefficiency issue, as they need to conduct element-wise linear recurrence in high dimension (i.e. $nd$). Since these element-wise operations cannot leverage tensor cores (a fast matrix multiply unit on GPUs), the dramatically increasing FLOPs and I/O costs would significantly slow down the training when $n$ is large. We notice that this is similar to the case in Mamba [^2] Though Mamba has an attention view (Ali et al., 2024) similar to that in linear attention, the attention computation cannot be written as matrix multiply like linear attention, and thus does not facilitate tensor core-based GPU acceleration., so Mamba needs a relatively small expansion ratio (i.e., $n=16$) and also a custom I/O efficient CUDA implementation to achieve a reasonably fast running speed.
In the next subsection, we explore another strategy that does not replace the dense projection matrices with structured ones, but changes the element-wise gating operations in Eq.1 to some other matrix/vector operations similar to linear attention, which allows for efficient training.
Table 3: PESE Ablation. We also provide the loss of Transformer (Xfmr), LLaMA (Xfmr++), and HGRN1 for references. $n$ standards for Expand Ratio. All experiments were trained on 10 billion tokens in a subset of the Pile.
Method $n$ Loss Params (B)
Xfmr - 2.362 0.380
Xfmr++ - 2.201 0.386
HGRN1 1 2.343 0.379
LR 4 2.246 0.385
8 2.247 0.386
GLT 4 2.333 0.386
GLTI 4 2.267 0.386
KRP 4 2.337 0.386
KP 4 2.332 0.386
HGRN2 4 2.255 0.385
8 2.237 0.385
128 2.195 0.385
3.2 HGRN2
The modification from HGRN1 to HGRN2 is simple yet effective. For the input gating mechanism, HGRN2 replaces the element-wise product with the outer product for state expansion. Consequently, $\mathbf{h}_{t}\in\mathbb{R}^{d\times d}$ and HGRN2 first diagonalize the forget gate vector and use the matrix dot product to update the hidden state. For the output gating mechanism, HGRN2 replaces the element-wise product with matrix-vector multiplication to project the expanded state to the original dimension size. The recurrent equation of HGRN2 is as follows,
$\mathbf{h}_{t}$ $=\mathrm{Diag}\{\mathbf{f}_{t}\}\cdot\mathbf{h}_{t-1}+(1-\mathbf{f}_{t})\otimes\mathbf{i}_{t}\in\mathbb{R}^{d\times d},$ (3)
$\mathbf{y_{t}}$ $=\mathbf{o}_{t}\cdot\mathbf{h}_{t}\in\mathbb{R}^{1\times d},$
where $\mathrm{Diag}$ stands for diagonalizing vectors, $\cdot$ is the matrix dot product, $\otimes$ is outer product. Note that this recurrent computation resembles that in GLA and RWKV-5/6.
Multihead variant.
Note that the complexity of recurrence increases dramatically from $O(BNd)$ to $O(BNd^{2})$ due to state expansion. Therefore, we introduce multi-head HGRN (similar to that in linear attention) such that the complexity is reduced to $O(BNd^{2}/H)$ for the number of heads $H$ and the state size effectively becomes $d^{2}/H$, i.e., the expansion ratio $n=d_{h}=d/H$ [^3] See Bolya et al. (2022) for more detailed complexity analysis. We conduct an ablation study on the expand ratio(head dimension) $n=\frac{d}{H}$ as shown in Figure 2. We can see that state expansion is very effective in improving language modeling performance. However, when the head dimension (i.e., state expansion ratio) is larger than 128, the performance gain is not too high. To balance computational cost and performance, we chose to use $d_{h}=128$ for the main experiments.
Figure 2: Expand Ratio (head dimension) Ablation. We tested the relationship between PPL (Perplexity) and expand ratio on the Wikitext-103 (Merity et al., 2017)(left) and a subset of the Pile (Gao et al., 2020) dataset(right). It can be seen that as the expand ratio increases, the PPL decreases.
Comparison to GLA.
Also note that this form of recurrence is similar to that of Gated Linear Attention (Yang et al., 2023) except for concrete parameterization. We list the correspondence between two parameterizations in Table 4. We can see that the output gate in HGRN2 amounts to the query in GLA, while the output gate in GLA is removed in HGRN2. The key vector in GLA corresponds to the input gate in HGRN2, which is tied to the forget gate, thus saving parameters.
HGRN2 GLA
$\mathbf{o}$ (output gate) $\mathbf{q}$ (query vector)
$\mathbf{1-f}$ (input gate) $\mathbf{k}$ (key vector)
$\mathbf{i}$ (input vector) $\mathbf{v}$ (value vector)
$\mathbf{f}$ (forget gate) $\bm{\alpha}$ (forget gate)
-- $\mathbf{o}$ (output gate)
Table 4: The correspondence between HGRN2 and GLA. Compared to GLA, HGRN2 removes GLA’s output gate and ties key vector to the forget gate (in terms of GLA terminology), thus saving parameters.
Hardware-efficient training.
Thanks to the similar computation structure to GLA, we can directly leverage their chunkwise algorithm and their highly optimized kernels  [^4] https://github.com/sustcsonglin/flash-linear-attention for hardware-efficient large-scale training. We refer the reader to their paper for more details.
Ending remarks.
Although HGRN2 shares many similarities with GLA, we believe that HGRN2 offers a unique perspective that is different from linear attention, starting from the approach of gated linear RNN. For example, it is not immediately apparent from the perspective of linear attention why key vectors should be constrained within the range of (0, 1) such as in Schlag et al. (2021) ; and why key vector and forget gate value should sum up to one. It becomes quite intuitive when one starts with the gated linear RNN framework and aims to explore state expansion.
4 Experiments
4.1 MQAR
Setting.
Multi-Query Associative Recall (MQAR) (Arora et al., 2023) is an enhanced version of the synthetic induction head dataset (Fu et al., 2023) to test the in-context associative recall ability for subquadractic models. Arora et al. (2023) found strong correlations between the accuracy of MQAR and the performance of language modeling. Our experimental setting strictly follows the original paper  [^5] https://github.com/HazyResearch/zoology, other technical details can also be found in the paper. Our hyperparameter sweeping range is: expand ratio $\in\{64,128\}$, learning rate $\in\{1e-5,5e-5,1e-4,5e-4,1e-3,5e-3,1e-2\}$.
Result.
We can see from Fig.3 that HGRN2 significantly outperforms HGRN1 across various model dimensions, showing the benefit of using state expansion.
Figure 3: Results on MQAR, where the x-axis represents the model dimension and the y-axis represents accuracy, with the task becoming more difficult as the sequence length increases. HGRN2 outperforms HGRN1 in all scenarios.
4.2 Language modeling
Table 5: Results on Wikitext-103. $\downarrow$ means lower is better. We adopted the configuration of HGRN1 44m for wikitext, and it can be seen that HGRN2 significantly outperforms HGRN1 as well as all other methods.
Model PPL (val)$\downarrow$ PPL (test)$\downarrow$ Params (M)
Attn-based
Transformer 24.40 24.78 44.65
FLASH 25.92 26.70 42.17
1+elu 27.44 28.05 44.65
Performer 62.50 63.16 44.65
cosFormer 26.53 27.06 44.65
MLP-based
Syn(D) 31.31 32.43 46.75
Syn(R) 33.68 34.78 44.65
gMLP 28.08 29.13 47.83
RNN-based
S4 38.34 39.66 45.69
DSS 39.39 41.07 45.73
GSS 29.61 30.74 43.84
RWKV-4 24.31 25.07 46.23
LRU 29.86 31.12 46.24
HGRN1 24.14 24.82 46.25
FFT-based
TNN 23.98 24.67 48.68
Ours
HGRN2 23.10 23.73 44.66
4.2.1 Wiki103
Setting.
For the Wikitext-103 experiment, we followed the configuration of HGRN1 to validate the performance of 44M models against a wide range of subquadractic models: FLASH (Hua et al., 2022), 1+elu (Katharopoulos et al., 2020), Performer (Choromanski et al., 2021), cosFormer (Qin et al., 2022b), Syn(D), Syn(R) (Tay et al., 2021a), gMLP (Liu et al., 2021), S4 (Gu et al., 2022a), DSS (Gupta et al., 2022b), RWKV-v4 (Peng et al., 2023), LRU (Orvieto et al., 2023), HGRN1 (Qin et al., 2023c), TNN (Qin et al., 2023a).
Result.
Table 5 shows the result. We can see that HGRN2 achieves the lowest perplexity among all compared subquadractic models. In particular, HGRN2 clearly outperforms HGRN1 using even fewer parameters.
Figure 4: We compared the loss curves of HGRN2, LLaMA, Mamba, and HGRN1 under the consumption of 30 billion tokens. HGRN2 is comparable to the state-of-the-art (SOTA) methods at 1 billion and better than SOTA methods at 3 billion size.
4.2.2 The Pile
Setting.
We conduct a controlled experiment in a relatively large-scale setting on the Pile (Gao et al., 2020) to compare the LLaMa architecture (i.e., Transformer++), Mamba (Gu & Dao, 2023), HGRN1 and HGRN2. Due to the limiation of computational resources, we only train the 1B/3B models for 30B tokens.
Result.
The training curve is shown in Fig. 4. We can see that, for 1B models, HGRN2 slightly underperforms Mamba. For 3B models, HGRN2 outperforms both LLaMa and Mamba, indicating the potential to further scale up HGRN2.
4.2.3 Downstream evaluation
We train 150M/350M/1B/3B HGRN2 for 100B tokens sampled from subsets within the Pile, C4 and Wikipedia datasets and evaluate them in Commonsense reasoning tasks. Table 6 shows the result. HGRN2 outperforms HGRN1 in almost all benchmarks and competes strongly against other models.
Table 6: Performance Comparison on Commonsense Reasoning Tasks. PS, T, HS, WG stand for parameter size (billion), tokens (billion), HellaSwag, and WinoGrande, respectively. $\dagger$ indicates the results reported by Yang et al. (2023)
Model Params Token PIQA HS WG ARC-e ARC-c OBQA AVG
GPT-Neo 0.13 300 63.06 30.40 50.43 43.73 23.12 26.20 39.49
OPT 0.16 300 62.95 31.35 50.43 43.52 22.70 28.00 39.83
Pythia 0.16 300 61.32 30.16 51.93 43.18 23.12 26.80 39.42
RWKV-4 0.17 - 65.07 32.26 50.83 47.47 24.15 29.60 41.56
HGRN1 0.15 100 65.02 33.33 50.20 46.68 23.81 28.60 41.27
HGRN2 0.15 100 66.43 35.44 51.70 46.63 24.32 28.40 42.15
OPT 0.35 300 64.58 36.69 52.49 44.02 23.89 28.20 41.65
Pythia 0.40 300 67.08 40.52 53.59 51.81 24.15 29.40 44.43
BLOOM 0.56 350 64.09 36.97 52.80 47.35 23.98 28.20 42.23
RWKV-4 0.43 - 67.52 40.90 51.14 52.86 25.17 32.40 45.00
HGRN1 0.35 100 66.70 38.12 51.70 49.20 25.26 30.60 43.60
HGRN2 0.39 100 69.97 46.16 52.72 53.58 23.98 32.40 46.47
GPT-Neo 1.3 300 71.11 48.93 54.93 56.19 25.85 33.60 48.44
OPT 1.3 300 71.71 53.70 59.35 57.24 29.69 33.20 50.82
Pythia 1.4 300 70.67 47.18 53.51 56.99 26.88 31.40 47.77
BLOOM 1.1 350 67.14 42.98 54.93 51.47 25.68 29.40 45.27
RWKV-4 1.5 - 72.36 52.48 54.62 60.48 29.44 34.00 50.56
${}^{\dagger}$ 1.3 100 71.90 49.90 54.40 61.80 28.40 34.20 50.10
${}^{\dagger}$ 1.3 100 71.40 51.20 54.10 60.10 30.40 33.80 50.20
${}^{\dagger}$ 1.3 100 69.80 47.50 51.10 58.50 27.40 31.80 47.68
HGRN1 1.0 100 70.89 48.02 51.62 55.64 27.90 31.60 47.61
HGRN2 1.0 100 74.16 54.85 56.12 58.71 27.22 34.00 50.84
OPT 2.7 300 73.83 60.60 61.01 60.77 31.31 35.2.0 53.79
Pythia 2.8 300 74.10 59.31 59.91 64.14 33.02 35.60 54.35
BLOOM 3.0 350 70.57 54.53 58.48 59.43 30.38 32.20 50.93
RWKV-4 3.0 - 72.42 58.75 57.3 62.92 35.15 36.20 53.79
HGRN2 2.9 100 74.59 59.92 57.46 62.58 29.86 35.00 53.24
Table 7: Performances Comparison on the Long Range Arena benchmark. The results of other models come from HGRN2 outperforms HGRN1 on all datasets. ${}^{\dagger}$ indicates the results reported by Alonso et al. (2024).
Model ListOps Text Retrieval Image Pathfinder Path-X AVG.
Transformer 38.37 61.95 80.69 40.57 65.26 - 47.81
cosFormer 36.50 67.70 83.15 51.23 71.96 - 51.76
FLASH 38.70 64.10 86.10 47.40 70.25 - 51.09
S4 59.60 86.82 90.90 88.65 94.20 96.35 86.09
DSS_softmax 60.60 84.80 87.80 85.70 84.60 87.80 81.88
DSSEXP 59.70 84.60 87.60 84.90 84.70 85.60 81.18
DSSEXP-NO-SCALE 59.30 82.40 86.00 81.20 81.30 - 65.03
TNN 61.04 87.90 90.97 88.24 93.00 96.10 86.21
S5 62.15 89.31 91.4 88.00 95.33 98.56 87.46
Mega 63.14 90.43 91.25 90.44 96.01 97.98 88.21
SGConv 61.45 89.2 91.11 87.97 95.46 97.83 87.17
LRU 60.20 89.40 89.90 89.00 95.10 94.20 86.30
${}^{\dagger}$ 38.02 82.98 72.14 69.82 69.26 67.32 66.59
${}^{\dagger}$ 32.34 71.75 66.58 61.15 73.38 69.53 62.45
HGRN1 59.95 88.14 94.23 88.69 92.92 97.50 86.91
HGRN2 60.52 88.97 95.07 89.33 93.95 98.12 87.66
4.3 LRA
Setting.
Long Range Arena (Tay et al., 2021b) is a benchmark for accessing the model’s long-dependency modeling ability. We use HGRN1’s setting, and compare wth existing methods shown below.
Baselines.
Transformer (Tay et al., 2021b), Cosformer (Qin et al., 2022c), FLASH (Hua et al., 2022), S4 (Gu et al., 2022b), DSS (Gupta et al., 2022a), TNN (Qin et al., 2023a), S5 (Smith et al., 2023), Mega (Ma et al., 2022), SGConv (Li et al., 2022), LRU (Orvieto et al., 2023), Mamba (Gu & Dao, 2023), Griffin (De et al., 2024).
Result.
Table 7 shows the result. We can see that HGRN2 outperforms HGRN1 and is competitive with other state-of-the-art models, while Mamba and Griffin failed to have a high accuracy, indicating the importance of lower bound as ablated in HGRN1.
4.4 Image modeling.
Setting
For the image classification task, we referred to the configuration of HGRN1 and train it on ImageNet-1k, also compared to TNN and the vanilla transformer
Result.
Table 8 shows the result. We can see that HGRN2 outperforms HGRN1 with a similar parameter size, while having advantage over previous TNN (Qin et al., 2023a) and Deit models (Touvron et al., 2021).
Table 8: Performances comparison of image classification on ImageNet-1k. HGRN2 performs favorably than competing methods with similar parameter sizes.
DeiT-Tiny DeiT-Small
Model Top1 Acc Param (M) Top1 Acc Parma (M)
Deit 72.20 5.7 79.90 22.0
TNN 72.29 6.4 79.20 23.4
HGRN1 74.40 6.1 80.09 23.7
HGRN2 75.39 6.1 80.12 23.8
5 Related work
Linear recurrent models.
Linear recurrent models mainly include linear RNNs, state-space models, and linear attention. State-space models (SSMs) are gaining great attention since the seminal work S4 (Gu et al., 2022a) and its more efficient diagonalized version (Gu et al., 2022c). Despite excellent performance in the LRA benchmark, it has been shown to have inferior performance in language modeling. Gating mechanisms have been shown to be crucial in improving SSMs’ language modeling performance (Mehta et al., 2023; Wang et al., 2022; Gu & Dao, 2023). Gupta et al. (2022c) build the connection between SSM and linear RNN. Orvieto et al. (2023) proposes a linear RNN layer (i.e., LRU) inspired by SSMs. Peng et al. (2023) successfully scale linear RNN models to billions of parameters for the first time.
For linear attention models, their language modeling performance has been underperforming softmax attention for a long time. Several improvements have been proposed to bridge the performance gap: (i) incorporating the forgetting mechanism (Peng et al., 2021; Schlag et al., 2021; Sun et al., 2023a; Qin et al., 2023b; Yang et al., 2023), (ii) using local attention (Qin et al., 2022a; Zhang et al., 2023; Arora et al., 2024), (iii) using higher-order polynomial feature map (Arora et al., 2024; Kacham et al., 2023) to make the resulting attention distribution more sharp (Zhang et al., 2024).
Gated linear recurrence.
Martin & Cundy (2018) first proposed a minimal gated linear recurrent layer and showed how to use the parallel scan algorithm to train linear RNNs in sequence-level parallel. Qin et al. (2023c) is largely based on this work with several adaptations and highlights the importance of data-dependent decay. De et al. (2024) build their model on LRU (Orvieto et al., 2023) and replace data-independent decays with data-dependent ones. They further use sliding-window attention to boost the performance. These models are limited in recurrent state size.
Gated recurrent models with matrix-valued recurrent state have been investigated in the literature of Neural Turing Machine (NTM, Graves et al. 2014) and linear Transformer (Katharopoulos et al., 2020). In NTM, the number of memory slots can be regarded as the state expansion ratio discussed in this work. NTM also included data-dependent decays in the form of erase vectors. However, NTM is hard to parallelize and thus slow to train in practice. ABC (Peng et al., 2022) could be considered a simplified and parallelizable version of NTM. The linear transformer is known to have the recurrent form (Katharopoulos et al., 2020) and is known to be closely related to fast weight programming (FWP Schlag et al. 2021). Gated FWPs have been investigated since Schlag & Schmidhuber (2017); Zhang & Zhou (2017), and have recently been revisited in Peng et al. (2021); Mao (2022); Yang et al. (2023); Katsch (2023); Pramanik et al. (2023). In particular, Yang et al. (2023) proposed a hardware-efficient training algorithm for these types of models.
6 Conclusion
In this work, we proposed HGRN2, which improves HGRN (Qin et al., 2023c) using an outer product-based state expansion mechanism inspired by linear attention, which allows for hardware-efficient training. Experiments on multiple tasks validate the advantages of HGRN2 over HGRN1. Large-scale language modeling experiments show that HGRN2 is competitive with other state-of-the-art models.
Appendix A Appendix
A.1 Experiment Configurations
In Table 9, the experiment configurations provided detail setups for both Auto-regressive Language Modeling (ALM) and ImageNet (IM) experiments, focusing on the WikiText-103 and ImageNet-1k datasets, respectively. ALM experiments utilize Byte Pair Encoding (BPE) with a vocabulary size of $50,265$ and sequence length of $512$, featuring a total batch size of $128$ and $50,000$ updates. ImageNet experiments differentiate between 6 million and 23 million parameter models, with total batch sizes of $1024$ and $2048$, both running for $300$ epochs but with differing warm-up periods. Optimization strategies vary between Adam for ALM and AdamW for IM, with specific learning rate schedulers and hyper-parameters tailored to each model’s scale. Additional configurations outline variations in model complexity, from $0.15$ to $2.9$ million parameters, adjusting layers, hidden dimensions, and GPUs used, aiming to comprehensively explore model performance across scales and setups.
Table 9: Comprehensive Configurations of the Model and Training Procedures for HGRN2 Experiments “Total batch size” means $\mathrm{batch\_per\_gpu}\times\mathrm{update\_freq}\times\mathrm{num\_gpus}$; “ALM” stands for Autoregressive Language Model; “IM” stands for Image Modeling.
ALM IM(6M) IM(23M)
Dataset WikiText-103 ImageNet-1k ImageNet-1k
Tokenizer method BPE - -
Src Vocab size 50265 - -
Sequence length 512 - -
Total batch size 128 1024 2048
Number of updates/epochs 50k updates 300 epochs 300 epochs
Warmup steps/epochs 4k steps 20 epochs 10 epochs
Peak learning rate 5e-4 7.5e-4 7.5e-4
Learning rate scheduler Inverse sqrt Cosine Cosine
Optimizer Adam Adamw Adamw
Adam $\epsilon$ 1e-8 1e-8 1e-8
Adam $(\beta_{1},\beta_{2})$ (0.9, 0.98) (0.9, 0.98) (0.9, 0.98)
Weight decay 0.1 0.05 0.1
Gradient clipping - 5.0 5.0
Table 10: Model configurations
Params Layers Hidden Dim Exp. Ratio L. R. Batch Size SeqLen GPUS
0.15 15 768 128 3.00E-04 26 2048 8
0.385 26 1024 128 3.00E-04 15 2048 8
1 18 2048 128 3.00E-04 10 2048 16
2.9 36 2560 128 3.00E-04 36 2048 64
A.2 Loss curve of HGRN2
The training loss curves for the HGRN2 models of different sizes—150M, 385M, and 1B, as shown in Fig. 5, which as the number of parameters increases, the model’s performance improves, with the 1B model consistently outperforming the others.
Figure 5: Training loss over train tokens for the 150m, 385m, 1B models. All models consumed approximately 100 billion tokens, and it can be observed that as the model size increases, the loss significantly decreases.

### v2

@@
 Hierarchically gated linear RNN (HGRN, Qin et al.
-2023c) has demonstrated competitive training speed and performance in language modeling, while offering efficient inference.
-However, the recurrent state size of HGRN remains relatively small, which limits its expressiveness.
-To address this issue, inspired by linear attention, we introduce a simple outer-product-based state expansion mechanism so that the recurrent state size could be significantly enlarged without introducing any additional parameters.
-The linear attention form also allows for hardware-efficient training.
-Our extensive experiments verify the advantage of HGRN2 over HGRN1 in language modeling, image classification, and Long Range Arena.
-Our largest 3B HGRN2 model slightly outperforms Mamba and LLaMa Architecture Transformer for language modeling in a controlled experiment setting; and performs competitively with many open-source 3B models in downstream evaluation but using much fewer total training tokens.
+2023c) has demonstrated competitive training speed and performance in language modeling while offering efficient inference.
+However, the recurrent state size of HGRN remains relatively small, limiting its expressiveness.
+To address this issue, we introduce a simple outer product-based state expansion mechanism, which significantly enlarges the recurrent state size without introducing any additional parameters.
+This enhancement also provides a linear attention interpretation for HGRN2, enabling hardware-efficient training.
+Our extensive experiments verify the advantage of HGRN2 over HGRN consistently across different settings and competitive with other recurrent models.
 1 Introduction
-Large language models (LLMs) have achieved great empirical success in recent years.
-However, serving Transformer-based LLMs is expensive due to the expensive KV cache management.
-Recurrent neural networks (RNNs), on the other hand, offer linear inference complexity with constant state size, which is ideal for serving.
-Therefore, there is great interest in studying parallelizable linear recurrent models such as linear RNNs (Peng et al., 2023; Orvieto et al., 2023; Qin et al., 2023c; De et al., 2024), linear attention (Sun et al., 2023b; Qin et al., 2023b; Yang et al., 2023; Arora et al., 2024; Peng et al., 2024), and state space models (Gu et al., 2022a; Smith et al., 2023; Gu & Dao, 2023).
+Large language models (LLMs) have achieved significant empirical success in recent years.
+However, serving Transformer-based LLMs is costly due to the expensive KV cache management.
+Recurrent neural networks (RNNs), on the other hand, offer linear inference complexity with constant state size, making them ideal for serving.
+Consequently, there is substantial interest in studying parallelizable linear recurrent models, such as linear RNNs (Peng et al., 2023; Orvieto et al., 2023; Qin et al., 2023c; De et al., 2024), linear attention (Sun et al., 2023; Qin et al., 2023b; Yang et al., 2023; 2024; Arora et al., 2024), and state space models (Gu et al., 2022a; Smith et al., 2023; Gu & Dao, 2023; Dao & Gu, 2024).
 RNNs have a fixed recurrent state size to encode all historical information.
-Therefore, it is important for RNNs (i) to better exploit the fix-sized states and (ii) to increase the recurrent state size.
-Recent improvements in linear RNN are exactly on this track, including the use of data-dependent decays and state expansion techniques.
-Data-dependent decays (also known as forget gates) are of great importance for RNNs (van der Westhuizen & Lasenby, 2018), which allows them to selectively retain useful information while erasing irrelevant information, so that the fixed-size recurrent state could be better exploited.
-HGRN (Qin et al., 2023c) first highlighted the importance of data-dependent decays for linear RNNs.
-Since then, many recent linear recurrent models have also employed data-dependent decays, such as Mamba (Gu & Dao, 2023), Gated Linear Attention (GLA, Yang et al.
-2023), Griffin (De et al., 2024), and RWKV-6 (Peng et al., 2024).
-However, HGRN did not increase the recurrent state size size, greatly restricted by the limited memory capacity, which prevents it from achieving LLaMa-like (Touvron et al., 2023a; b) language modeling performance, as noted in Qin et al.
+Therefore, it is important for RNNs to (i) utilize the fixed-sized states effectively and (ii) increase the recurrent state size to enhance memory capacity.
+Recent improvements in linear RNNs follow this approach, incorporating techniques such as data-dependent decays and state expansion.
+Data-dependent decays (also known as forget gates) are crucial for RNNs (van der Westhuizen & Lasenby, 2018), allowing them to selectively retain useful information while erasing irrelevant information.
+This enables the fixed-size recurrent state to store only important information more efficiently.
+HGRN (Qin et al., 2023c) first emphasized the importance of data-dependent decays for linear RNNs.
+Many recent linear recurrent models, such as Mamba (Gu & Dao, 2023), Gated Linear Attention (GLA, Yang et al.
+2023), Griffin (De et al., 2024), and RWKV-6 (Peng et al., 2024), also employ data-dependent decays.
+However, HGRN did not increase the recurrent state size, which is greatly restricted by limited memory capacity.
+This limitation prevents it from achieving LLaMa-like (Touvron et al., 2023a; b) language modeling performance, as noted in Qin et al.
 (2024).
-Recent state-of-the-art linear recurrent models, such as Mamba, GLA, and RWKV-6, additionally use state expansion techniques, which greatly increase the recurrent state size and thereby enhance the memory capacity.
-In this work, we propose HGRN2, with the aim of increasing the size of the recurrent state for HGRN while retaining both parameter efficiency and training efficiency.
-We first explore structured matrices to directly expand the state size in a parameter-efficient manner.
-Empirically, we found that they improved the language modeling performance, but still faced with the training inefficiency issue, thus they cannot scale the recurrent state size well.
-Inspired by linear attention, we then explore using a nonparameteric outer product-based state expansion mechanism, which allows efficient scaling of the size of the recurrent state during training without introducing any additional parameters.
-Thanks to the gated linear attention form, we could borrow the hardware-efficient GLA training algorithm described in Yang et al.
+Recent state-of-the-art linear recurrent models, such as Mamba, GLA, and RWKV-6, have addressed this issue by employing state-expansion techniques.
+These techniques significantly increase the recurrent state size and thereby enhance memory capacity, which has been shown to be crucial for language modeling performance and directly correlated with retrieval ability (Arora et al., 2024).
+In this work, we propose HGRN2, which aims to increase the recurrent state size for HGRN while retaining both parameter and training efficiency.
+We first explore structured matrices to expand the state size directly in a parameter-efficient manner.
+Empirically, we found that this approach improves language modeling performance but still encounters training inefficiencies, which limit the scaling of the recurrent state size.
+Inspired by linear attention, we then explore using a non-parametric outer product-based state expansion mechanism.
+This approach allows for efficient scaling of the recurrent state size during training without introducing additional parameters.
+Due to the matrix multiply form of linear attention, we can leverage the hardware-efficient linear attention training algorithm described in Yang et al.
+(2023); Qin et al.
 (2024) for large-scale experiments.
-We extensively evaluate HGRN2 in language modeling, image classification, and the Long Range Arena benchmark.
-HGRN2 outperforms HGRN1 in all these benchmarks, showing the significant benefit of using state expansion.
-Our largest 3B HGRN2 model slightly outperforms Mamba and LLaMa Architecture Transformer for language modeling in a controlled experiment setting; and performs competitively with many open-source 3B models in downstream evaluation but using much fewer total training tokens.
+As a result, HGRN2 can be regarded as an improved parameterization of GLA.
+We extensively evaluate HGRN2 across various tasks, demonstrating that it consistently outperforms HGRN1 in multiple domains.
+In language modeling, we show HGRN2 to be highly competitive compared to other subquadratic efficient models.
 2 Background
 2.1 Gated linear RNN
-Given input $\mathbf{x}\in\mathbb{R}^{N\times d}$ where the sequence length is $N$ and the model dimension is $d$.
-A minimalist gated linear recurrent layer transforms the input $\mathbf{x}$ into hidden states $\mathbf{h}\in\mathbb{R}^{N\times d}$ and the output $\mathbf{y}\in\mathbb{R}^{N\times d}$, as defined below.
+Given input $\mathbf{x}\in\mathbb{R}^{N\times d}$, where the sequence length is $N$ and the model dimension is $d$, a minimalist gated linear recurrent layer (Martin & Cundy, 2018) transforms the input $\mathbf{x}$ into hidden states $\mathbf{h}\in\mathbb{R}^{N\times d}$ and the output $\mathbf{y}\in\mathbb{R}^{N\times d}$, as defined below:
 $\mathbf{g}_{t}$ $=\sigma\left(\mathbf{U}\mathbf{x}_{t}+\mathbf{b}_{u}\right),$ (1)
-$\mathbf{i}_{t}$ $=\mathbf{V}\mathbf{x}_{t}+\mathbf{b}_{v},$
-$\mathbf{o}_{t}$ $=\tau\left(\mathbf{W}\mathbf{x}_{t}+\mathbf{b}_{w}\right),$
+$\mathbf{i}_{t}$ $=\tau\left(\mathbf{V}\mathbf{x}_{t}+\mathbf{b}_{v}\right),$
+$\mathbf{o}_{t}$ $=\sigma\left(\mathbf{W}\mathbf{x}_{t}+\mathbf{b}_{w}\right),$
 $\mathbf{h}_{t}$ $=\mathbf{g}_{t}\odot\mathbf{h}_{t-1}+\left(1-\mathbf{g}_{t}\right)\odot\mathbf{i}_{t},$
 $\mathbf{y}_{t}$ $=\mathbf{h}_{t}\odot\mathbf{o}_{t},$
-where $\odot$ is element-wise product; $\sigma$ is the sigmoid function and $\tau$ is nonlinear activation (we choose to use $\mathrm{SiLU}$); $\mathbf{i}_{t}$ is the input vector; $\mathbf{g}_{t},\mathbf{o}_{t}$ are the forget gate and output gate, respectively.
-The input gate is tied to the forget gate: $1-\mathbf{g}_{t}$, which is commonly used in many gated RNNs such as GRU (Chung et al., 2014).
+where $\odot$ denotes element-wise product; $\sigma$ is the sigmoid function, and $\tau$ is a nonlinear activation function (we choose to use $\mathrm{SiLU}$); $\mathbf{i}_{t}$ is the input vector; $\mathbf{g}_{t}$ and $\mathbf{o}_{t}$ are the forget gate and output gate, respectively.
+The input gate is tied to the forget gate as $1-\mathbf{g}_{t}$, a common approach used in many gated RNNs such as GRU (Chung et al., 2014).
 2.2 HGRN (Qin et al., 2023c)
-Compared to Eq. 1, HGRN makes two adjustments: (i) complex-valued recurrence. (ii) forget gates with monotonically increased lower bound values from bottom layers to upper layers.
-For (i), similarly to the findings of Gu & Dao (2023) and De et al.
-(2024), we empirically found that the complex-valued recurrence is not necessary, as shown in Table 1.
-We speculate that the reason Qin et al.
-(2023c) found it useful is due to state expansion: The complex-valued recurrent state is twice the size of that in the real-valued recurrent state.
-If we directly expand the real-valued recurrent state size from $d$ to $2d$, the language modeling performance on Wikitext-103 corpus is even better.
+Compared to Eq. 1, HGRN makes two adjustments: (i) complex-valued recurrence and (ii) forget gates with monotonically increased lower bound values from bottom layers to upper layers.
+For (i), similar to the findings in Gu & Dao (2023) and De et al.
+(2024), we empirically found that complex-valued recurrence is not necessary, as shown in Table 1.
+The reason why HGRN found it useful is due to state expansion: the complex-valued recurrent state is twice the size of that in the real-valued recurrent state.
+If we directly expand the real-valued recurrent state size from $d$ to $2d$, the language modeling performance on the Wikitext-103 corpus is even better.
 Therefore, we only consider the real-valued recurrence thereafter.
@@
 Real HGRN1 $d$ 25.34 26.12 46.24
-Real HGRN1 $2d$ 23.81 24.55 47.82
-For (ii), suppose that the total number of layers is $L$, and HGRN introduces a data-independent learnable matrix $\Gamma\in\mathbb{R}^{L\times d}$.
-$\Gamma_{i}$ will be the lowest values of the forget gate for the $i$th layer in all time steps.
+Real HGRN1 $2d$ 24.04 24.64 45.46
+For (ii), suppose the total number of layers is $L$.
+HGRN introduces a data-independent learnable matrix $\Gamma\in\mathbb{R}^{L\times d}$, where $\Gamma_{i}$ represents the lowest values of the forget gate for the $i$-th layer at all time steps.
 HGRN argues that this lower bound should be monotonically increasing from bottom to top, encouraging the bottom layers to model short-term local dependencies and the upper layers to model long-term dependencies.
-To enforce such monotonicity, HGRN used the cumulative softmax operator cumax [^1] This was first introduced by Shen et al.
-(2018):
-$\beta:=\texttt{cumax}(\Gamma)=\texttt{cumsum}(\texttt{softmax}(\Gamma,\mathrm{dim=0}),\mathrm{dim=0})\in\mathbb{R}^{L\times d},\beta^{i}=[\beta]_{i}\in\mathbb{R}^{d}.$
-To avoid the lower bound being one in the highest layer, HGRN subtracts all $\beta$ by $\beta^{0}$, so that the lower bound for the first layer is zero.
-After obtaining the lower bound values, the old forget gate $\mathbf{g}_{t}$ learns the residuals instead, resulting the new forget gate $\mathbf{f}_{t}$
+To enforce this monotonicity, HGRN uses the cumulative softmax operator cumax (Shen et al., 2018):
+$\beta:=\texttt{cumax}(\Gamma)=\texttt{cumsum}(\texttt{softmax}(\Gamma,\text{dim}=0),\text{dim}=0)\in\mathbb{R}^{L\times d},\quad\beta^{i}=[\beta]_{i}\in\mathbb{R}^{d}.$
+To prevent the lower bound from reaching one in the highest layer, HGRN subtracts all $\beta$ values by $\beta^{0}$, making the lower bound for the first layer zero.
+After obtaining the lower bound values, the forget gate $\mathbf{g}_{t}$ learns residuals instead, resulting in a new forget gate $\mathbf{f}_{t}$:
 $\mathbf{f}_{t}^{i}$ $=\beta^{i}+(1-\beta^{i})\odot\mathbf{g}_{t}^{i},$ (2)
 $\mathbf{h}_{t}^{i}$ $=\mathbf{f}_{t}^{i}\odot\mathbf{h}_{t-1}^{i}+(1-\mathbf{f}_{t}^{i})\odot\mathbf{i}_{t}^{i},$
-where we use superscript to indicate the layer index.
-Figure 1: The neural architecture of HGRN2.
-Each HGRN2 layer includes a token mixer layer HGRU2 and a channel mixerlayer GLU.
-HGRU2 employs recurrent computation through Eq.
-3, where $\mathbf{i}_{t}$ is the input vector, $\mathbf{g}_{t}$ is the forget gate (not lower bounded), $\beta^{i}$ is the lower bound of the forget gate value, $\mathbf{o}_{t}$ is the output gate for layer $i$.
+where the superscript indicates the layer index. This additive lower bound approach has been shown to mitigate the issue of saturated gates (Gu et al., 2020).
 3 Method
 3.1 Explorations of state expansion methods
-The goal is to scale the size of the HGRN recurrent state from $d$ to $nd$, where $n$ is the expansion ratio of the state.
+The goal of this work is to scale the size of the HGRN recurrent state from $d$ to $nd$, where $n$ is the state expansion ratio.
 However, if we use the original parameterization in Eq.
-1, $\mathbf{U},\mathbf{V},\mathbf{W}$ will be of size $d\times nd$, which is very parameter inefficient when $n$ is large.
-Ideally, the number of parameters should be around $d^{2}$ as in the original case for each projection.
-First, we consider the following structured matrices, including low-rank, group linear transformation (GLT, Mehta et al.
-2018; 2021), group linear transformation with group interaction (GLTI), Khatri-Rao product (KRP) and Kronecker product (KP), to replace the dense projection matrix $\mathbb{R}^{d}\rightarrow\mathbb{R}^{nd}$ as described in Table 2.
-Table 2: We use Einstein Summation notation to represent Parameter Efficient State Expand (PESE) methods.
-We use Blue to represent the input, Black to represent data independent weight and Red to represent the output.
-PESE Naive LR GLT
-Equation ${{\color[rgb]{0,0,1}d},d\ nd\to{\color[rgb]{1,0,0}nd}}$ ${\color[rgb]{0,0,1}d},d\ r,r\ nd\to{\color[rgb]{1,0,0}nd}$ $\color[rgb]{0,0,1}d=(n\ e)\to n\ e$ ${\color[rgb]{0,0,1}n\ e},n\ e\ d\to{\color[rgb]{1,0,0}(n\ d)}$
-Prameters $nd^{2}$ $dr(n+1)\approx d^{2}$ $d^{2}$
-PESE GLTI KRP KP
-Equation $\color[rgb]{0,0,1}d=(n\ e)\to n\ e$ ${\color[rgb]{0,0,1}n^{\prime}\ e},n^{\prime}\ e\ d,n^{\prime}\ n\to{\color[rgb]{1,0,0}(n\ d)}$ ${\color[rgb]{0,0,1}d},n\ d\to{\color[rgb]{1,0,0}(n\ d)}$ ${\color[rgb]{0,0,1}d},d\ d,n\to{\color[rgb]{1,0,0}(n\ d)}$
-Parameters $d^{2}+n^{2}$ $nd$ $d^{2}+n$
-After obtaining expanded $\mathbf{g},\mathbf{i},\mathbf{o}$, we feed them into element-wise gated linear recurrent layers as in Eq 1 and  2, resulting in the output vector $\mathbf{y}_{t}\in\mathbb{R}^{n\times d}$.
-To project the expanded dimension back to the original dimension, we simply marginalize the dimension corresponding to $n$.
-The result is shown in Table 3. We found that state expansion generally improves the performance and that the low-rank matrix performs the best among these candidates.
-However, these PESE methods face the training inefficiency issue, as they need to conduct element-wise linear recurrence in high dimension (i.e.
-$nd$).
-Since these element-wise operations cannot leverage tensor cores (a fast matrix multiply unit on GPUs), the dramatically increasing FLOPs and I/O costs would significantly slow down the training when $n$ is large.
-We notice that this is similar to the case in Mamba [^2] Though Mamba has an attention view (Ali et al., 2024) similar to that in linear attention, the attention computation cannot be written as matrix multiply like linear attention, and thus does not facilitate tensor core-based GPU acceleration., so Mamba needs a relatively small expansion ratio (i.e., $n=16$) and also a custom I/O efficient CUDA implementation to achieve a reasonably fast running speed.
-In the next subsection, we explore another strategy that does not replace the dense projection matrices with structured ones, but changes the element-wise gating operations in Eq.1 to some other matrix/vector operations similar to linear attention, which allows for efficient training.
-Table 3: PESE Ablation.
-We also provide the loss of Transformer (Xfmr), LLaMA (Xfmr++), and HGRN1 for references.
-$n$ standards for Expand Ratio.
-All experiments were trained on 10 billion tokens in a subset of the Pile.
-Method $n$ Loss Params (B)
-Xfmr - 2.362 0.380
-Xfmr++ - 2.201 0.386
-HGRN1 1 2.343 0.379
-LR 4 2.246 0.385
-8 2.247 0.386
-GLT 4 2.333 0.386
-GLTI 4 2.267 0.386
-KRP 4 2.337 0.386
-KP 4 2.332 0.386
-HGRN2 4 2.255 0.385
-8 2.237 0.385
-128 2.195 0.385
+1, the matrices $\mathbf{U},\mathbf{V},\mathbf{W}$ will have dimensions $d\times nd$, which becomes very parameter inefficient when $n$ is large.
+Ideally, the number of parameters should be around $d^{2}$, as in the original case for each projection.
+To achieve this, we first consider using structured matrices (e.g., low-rank matrices) to replace the dense projection matrix $\mathbb{R}^{d}\rightarrow\mathbb{R}^{nd}$, as described in Table 2.
+Table 2: Parameter Efficient State Expansion (PESE) methods using Einstein Summation notation.
+Blue represents the input, Black represents data-independent weights, and Red represents the output.
+We list the Einstein Summation for low-rank (LR), group linear transformation (GLT), group linear transformation with interaction (GLTI), Khatri-Rao product (KRP), and Kronecker product (KP).
+      Method       Equation       Parameter #
+      Naive       ${{\color[rgb]{0,0,1}d},\mathbf{d\ nd}\rightarrow{\color[rgb]{1,0,0}nd}}$       $nd^{2}$
+      LR       ${\color[rgb]{0,0,1}d},\mathbf{d\ r,\ r\ nd}\rightarrow{\color[rgb]{1,0,0}nd}$       $dr(n+1)\approx d^{2}$
+      GLT              $\color[rgb]{0,0,1}d=(n\ e)\rightarrow n\ e$       ${\color[rgb]{0,0,1}n\ e},\mathbf{n\ e\ d}\rightarrow{\color[rgb]{1,0,0}n\ d}$       $d^{2}$
+      GLTI              $\color[rgb]{0,0,1}d=(n\ e)\rightarrow n\ e$       ${\color[rgb]{0,0,1}n\ e},\mathbf{n\ e\ d,\ n\ n}\rightarrow{\color[rgb]{1,0,0}nd}$       $d^{2}+n^{2}$
+      KRP       ${\color[rgb]{0,0,1}d},\mathbf{n\ d}\rightarrow{\color[rgb]{1,0,0}nd}$       $nd$
+      KP       ${\color[rgb]{0,0,1}d},\mathbf{d\ d,\ n}\rightarrow{\color[rgb]{1,0,0}nd}$       $d^{2}+n$
+After obtaining the expanded $\mathbf{g},\mathbf{i},\mathbf{o}$, we feed them into element-wise gated linear recurrent layers as in Eq.
+1 and Eq.
+2, resulting in the output vector $\mathbf{y}_{t}\in\mathbb{R}^{n\times d}$.
+To project the expanded dimension back to the original dimension, we simply sum over the dimension corresponding to $n$.
+The results are shown in Table 3. We found that state expansion generally improves performance, with the low-rank matrix performing the best among these candidates.
+Table 3: PESE Ablation. Ablation studies on various parameter-efficient methods, as described in Table 2. Each model was trained on 10 billion tokens from the Pile dataset.
+Method n PPL Params (M)
+Xfmr - 5.16 380
+Xfmr++ - 4.62 386
+HGRN1 1 5.10 379
+LR 4 4.76 385
+8 4.77 386
+GLT 4 5.06 386
+GLTI 4 4.83 386
+KRP 4 5.08 386
+KP 4 5.06 386
+HGRN2 4 4.79 385
+8 4.73 385
+128 4.62 385
+However, these methods face training inefficiency issues, as they require conducting element-wise linear recurrence in high dimensions (i.e., $nd$).
+Since these element-wise operations cannot leverage tensor cores (a fast matrix multiplication unit on GPUs), the dramatically increasing FLOPs and I/O costs significantly slow down training when $n$ is large.
+We notice that this is similar to the case in Mamba [^1] Though Mamba has an attention mechanism (Ali et al., 2024) similar to that in linear attention, the attention computation cannot be expressed as a matrix multiplication like linear attention, and thus does not facilitate tensor core-based GPU acceleration, as well acknowledged in Mamba2 (Dao & Gu, 2024).
+, which requires a relatively small expansion ratio (i.e., $n=16$) and a custom I/O-efficient CUDA implementation to achieve a reasonably fast running speed.
+In the next subsection, we explore an alternative strategy that does not replace the dense projection matrices with structured ones but instead changes the element-wise gating operations in Eq.1 to other matrix/vector operations similar to those used in linear attention.
+This approach allows for more efficient training.
 3.2 HGRN2
+Figure 1: Network Structure of HGRN2.
+Each HGRN2 layer includes a token mixer layer, HGRU2, and a channel mixer, GLU.
+HGRU2 employs recurrent computation as described in Eq.
+3, where $\mathbf{i}_{t}$ is the input state, $\mathbf{g}_{t}$ is the forget gate, $\mathbf{o}_{t}$ is the output gate, and $\beta^{i}$ is the lower bound for layer $i$.
 The modification from HGRN1 to HGRN2 is simple yet effective.
-For the input gating mechanism, HGRN2 replaces the element-wise product with the outer product for state expansion.
-Consequently, $\mathbf{h}_{t}\in\mathbb{R}^{d\times d}$ and HGRN2 first diagonalize the forget gate vector and use the matrix dot product to update the hidden state.
-For the output gating mechanism, HGRN2 replaces the element-wise product with matrix-vector multiplication to project the expanded state to the original dimension size.
-The recurrent equation of HGRN2 is as follows,
-$\mathbf{h}_{t}$ $=\mathrm{Diag}\{\mathbf{f}_{t}\}\cdot\mathbf{h}_{t-1}+(1-\mathbf{f}_{t})\otimes\mathbf{i}_{t}\in\mathbb{R}^{d\times d},$ (3)
-$\mathbf{y_{t}}$ $=\mathbf{o}_{t}\cdot\mathbf{h}_{t}\in\mathbb{R}^{1\times d},$
-where $\mathrm{Diag}$ stands for diagonalizing vectors, $\cdot$ is the matrix dot product, $\otimes$ is outer product. Note that this recurrent computation resembles that in GLA and RWKV-5/6.
+For the input gate, HGRN2 replaces the element-wise product with the outer product for state expansion.
+Consequently, $\mathbf{h}_{t}\in\mathbb{R}^{d\times d}$, and HGRN2 first diagonalizes the forget gate vector and uses the matrix dot product to update the hidden state.
+For the output gate, HGRN2 replaces the element-wise product with matrix-vector multiplication to project the expanded state back to the original dimension.
+The recurrent equation of HGRN2 is as follows:
+$\mathbf{h}_{t}$ $=\mathbf{h}_{t-1}\cdot\mathrm{Diag}\{\mathbf{f}_{t}\}+\mathbf{i}_{t}\otimes(1-\mathbf{f}_{t})\in\mathbb{R}^{d\times d},$ (3)
+$\mathbf{y}_{t}$ $=\mathbf{h}_{t}\cdot\mathbf{o}_{t}\in\mathbb{R}^{d},$
+where $\mathrm{Diag}$ denotes the diagonalization of vectors, $\cdot$ represents the matrix dot product, and $\otimes$ indicates the outer product.
 Multihead Variant.
-Note that the complexity of recurrence increases dramatically from $O(BNd)$ to $O(BNd^{2})$ due to state expansion.
-Therefore, we introduce multi-head HGRN (similar to that in linear attention) such that the complexity is reduced to $O(BNd^{2}/H)$ for the number of heads $H$ and the state size effectively becomes $d^{2}/H$, i.e., the expansion ratio $n=d_{h}=d/H$ [^3] See Bolya et al.
+The complexity of recurrence increases dramatically from $O(BNd)$ to $O(BNd^{2})$ due to state expansion.
+To address this, we introduce a multihead variant of HGRN (similar to that in linear attention) such that the complexity is reduced to $O(BNd^{2}/H)$ for the number of heads $H$, effectively making the state size $d^{2}/H$, i.e., the expansion ratio $n=d_{h}=d/H$ [^2] See Bolya et al.
 (2022) for more detailed complexity analysis.
-We conduct an ablation study on the expand ratio(head dimension) $n=\frac{d}{H}$ as shown in Figure 2.
-We can see that state expansion is very effective in improving language modeling performance.
-However, when the head dimension (i.e., state expansion ratio) is larger than 128, the performance gain is not too high.
-To balance computational cost and performance, we chose to use $d_{h}=128$ for the main experiments.
+We conducted an ablation study on the expansion ratio (or head dimension) $n=\frac{d}{H}$, as shown in Figure 2.
+The results show that state expansion significantly improves language modeling performance.
+However, when the head dimension (i.e., state expansion ratio) exceeds 128, the performance gain diminishes.
+To balance computational cost and performance, we chose $d_{h}=128$ for the main experiments.
 Figure 2: Expand Ratio (Head Dimension) Ablation.
-We tested the relationship between PPL (Perplexity) and expand ratio on the Wikitext-103 (Merity et al., 2017)(left) and a subset of the Pile (Gao et al., 2020) dataset(right).
-It can be seen that as the expand ratio increases, the PPL decreases.
+We tested the relationship between PPL (Perplexity) and the expand ratio on the Wikitext-103 (Merity et al., 2017) dataset (left) and a subset of the Pile (Gao et al., 2020) dataset (right).
 Comparison to GLA.
-Also note that this form of recurrence is similar to that of Gated Linear Attention (Yang et al., 2023) except for concrete parameterization.
-We list the correspondence between two parameterizations in Table 4.
-We can see that the output gate in HGRN2 amounts to the query in GLA, while the output gate in GLA is removed in HGRN2.
-The key vector in GLA corresponds to the input gate in HGRN2, which is tied to the forget gate, thus saving parameters.
+It is important to note that the recurrence form in HGRN2 is identical to that of GLA (Yang et al., 2023), except for the specific parameterization.
+We list the correspondences between the two parameterizations in Table 4.
+As shown, the output gate in HGRN2 corresponds to the query in GLA, while the output gate in GLA is omitted in HGRN2.
+The key vector in GLA corresponds to the input gate in HGRN2, which is tied to the forget gate, thereby saving parameters.
+Table 4: The correspondence between HGRN2 and GLA is as follows.
                  HGRN2                  GLA
@@
                  $\mathbf{f}$ (forget gate)                  $\bm{\alpha}$ (forget gate)
--- $\mathbf{o}$ (output gate)
-Table 4: The correspondence between HGRN2 and GLA. Compared to GLA, HGRN2 removes GLA’s output gate and ties key vector to the forget gate (in terms of GLA terminology), thus saving parameters.
+                 --                  $\mathbf{o}$ (output gate)
 Hardware-Efficient Training.
-Thanks to the similar computation structure to GLA, we can directly leverage their chunkwise algorithm and their highly optimized kernels  [^4] https://github.com/sustcsonglin/flash-linear-attention for hardware-efficient large-scale training.
-We refer the reader to their paper for more details.
-Ending remarks.
-Although HGRN2 shares many similarities with GLA, we believe that HGRN2 offers a unique perspective that is different from linear attention, starting from the approach of gated linear RNN.
-For example, it is not immediately apparent from the perspective of linear attention why key vectors should be constrained within the range of (0, 1) such as in Schlag et al.
-(2021) ; and why key vector and forget gate value should sum up to one.
-It becomes quite intuitive when one starts with the gated linear RNN framework and aims to explore state expansion.
+Due to its computational structure’s similarity to GLA, we can directly leverage their chunkwise algorithm and highly optimized kernels for hardware-efficient large-scale training.
+For more details, we refer readers to their paper.
+Concluding Remarks.
+Although HGRN2 shares many similarities with GLA, we believe that HGRN2 offers a unique perspective distinct from linear attention, originating from the approach of gated linear RNNs.
+For instance, it may not be immediately clear from the perspective of linear attention why key vectors should be constrained within the range of (0, 1) or why the key vector and forget gate value should sum to one.
+However, these concepts become quite intuitive when starting from the gated linear RNN framework and exploring state expansion.
 4 Experiments
@@
 Setting.
-Multi-Query Associative Recall (MQAR) (Arora et al., 2023) is an enhanced version of the synthetic induction head dataset (Fu et al., 2023) to test the in-context associative recall ability for subquadractic models.
+Multi-Query Associative Recall (MQAR) (Arora et al., 2023) is an enhanced version of the synthetic induction head dataset (Fu et al., 2023), designed to test the in-context associative recall ability of subquadratic models.
 Arora et al.
-(2023) found strong correlations between the accuracy of MQAR and the performance of language modeling.
-Our experimental setting strictly follows the original paper  [^5] https://github.com/HazyResearch/zoology, other technical details can also be found in the paper.
-Our hyperparameter sweeping range is: expand ratio $\in\{64,128\}$, learning rate $\in\{1e-5,5e-5,1e-4,5e-4,1e-3,5e-3,1e-2\}$.
+(2023) found strong correlations between MQAR accuracy and language modeling performance.
+Our experimental setting strictly follows the original paper [^3] https://github.com/HazyResearch/zoology.
+Our hyperparameter sweep included the following ranges: expansion ratio $\in\{64,128\}$ and learning rate $\in\{1e-5,5e-5,1e-4,5e-4,1e-3,5e-3,1e-2\}$.
 Result.
-We can see from Fig.3 that HGRN2 significantly outperforms HGRN1 across various model dimensions, showing the benefit of using state expansion.
-Figure 3: Results on MQAR, where the x-axis represents the model dimension and the y-axis represents accuracy, with the task becoming more difficult as the sequence length increases.
+As shown in Fig.
+3, HGRN2 significantly outperforms HGRN1 across various model dimensions, demonstrating the benefits of state expansion in improving memory capacity and, consequently, in-context recall ability.
+Figure 3: Results on MQAR, where the x-axis represents the model dimension and the y-axis represents accuracy.
+The task becomes more challenging as the sequence length increases.
 HGRN2 outperforms HGRN1 in all scenarios.
 4.2 Language modeling
+4.2.1 Wikitext-103
 Table 5: Results on Wikitext-103.
-$\downarrow$ means lower is better.
-We adopted the configuration of HGRN1 44m for wikitext, and it can be seen that HGRN2 significantly outperforms HGRN1 as well as all other methods.
-Model PPL (val)$\downarrow$ PPL (test)$\downarrow$ Params (M)
-Attn-based
+Model PPL (val) PPL (test) Params (M)
 Transformer 24.40 24.78 44.65
@@
 cosFormer 26.53 27.06 44.65
-MLP-based
 Syn(D) 31.31 32.43 46.75
@@
 gMLP 28.08 29.13 47.83
-RNN-based
 S4 38.34 39.66 45.69
@@
 LRU 29.86 31.12 46.24
+TNN 23.98 24.67 48.68
+Mamba 22.58 23.19 44.39
 HGRN1 24.14 24.82 46.25
-FFT-based
-TNN 23.98 24.67 48.68
-Ours
 HGRN2 23.10 23.73 44.66
-4.2.1 Wiki103
 Setting.
-For the Wikitext-103 experiment, we followed the configuration of HGRN1 to validate the performance of 44M models against a wide range of subquadractic models: FLASH (Hua et al., 2022), 1+elu (Katharopoulos et al., 2020), Performer (Choromanski et al., 2021), cosFormer (Qin et al., 2022b), Syn(D), Syn(R) (Tay et al., 2021a), gMLP (Liu et al., 2021), S4 (Gu et al., 2022a), DSS (Gupta et al., 2022b), RWKV-v4 (Peng et al., 2023), LRU (Orvieto et al., 2023), HGRN1 (Qin et al., 2023c), TNN (Qin et al., 2023a).
+For the Wikitext-103 experiment, we followed the configuration of HGRN1 to validate the performance of 44M models against a wide range of subquadratic models: FLASH (Hua et al., 2022), 1+elu (Katharopoulos et al., 2020), Performer (Choromanski et al., 2021), cosFormer (Qin et al., 2022b), Syn(D), Syn(R) (Tay et al., 2021a), gMLP (Liu et al., 2021), S4 (Gu et al., 2022a), DSS (Gupta & Berant, 2022), RWKV-v4 (Peng et al., 2023), LRU (Orvieto et al., 2023), HGRN1 (Qin et al., 2023c), TNN (Qin et al., 2023a), and Mamba (Gu & Dao, 2023).
+All reported results are from our own runs under the same settings.
 Result.
-Table 5 shows the result. We can see that HGRN2 achieves the lowest perplexity among all compared subquadractic models. In particular, HGRN2 clearly outperforms HGRN1 using even fewer parameters.
-Figure 4: We compared the loss curves of HGRN2, LLaMA, Mamba, and HGRN1 under the consumption of 30 billion tokens.
-HGRN2 is comparable to the state-of-the-art (SOTA) methods at 1 billion and better than SOTA methods at 3 billion size.
+Table 5 shows the results. HGRN2 clearly outperforms HGRN1 but slightly underperforms Mamba.
+4.2.2 Slimpajama
+We conducted language modeling experiments with 1.3B and 2.7B parameters on the Slimpajama dataset (Soboleva et al., 2023), using the FlashLinearAttention (Yang & Zhang, 2024) codebase for training.
+[^4] Model checkpoints are available at https://huggingface.co/fla-hub.
+The results, shown in Table 6, demonstrate that HGRN2 consistently outperforms other competitive linear recurrent models across three model scales.
+This suggests that HGRN2 provides a superior parameterization compared to GLA, as both models share an identical recurrent structure.
+Table 6: Slimpajama language modeling results.
+Lamb. Wiki. ARCe ARCc Hella. Lamb. PIQA Wino. Avg
+[…]
+1.3B parameters with 100B training tokens
+Transformer++ 15.3 17.1 54.1 27.1 49.3 47.0 70.3 54.9 50.5
+Mamba 16.5 18.2 57.3 26.6 48.1 43.4 69.5 53.7 49.8
+RetNet 15.4 17.3 57.4 27.9 50.3 44.6 71.7 51.8 50.6
+GLA 15.4 17.6 55.4 27.7 49.0 46.4 69.9 54.0 50.4
+HGRN2 11.8 16.9 58.1 28.1 51.8 49.4 71.4 52.3 51.9
+2.7B parameters with 100B training tokens
+Transformer++ 10.7 15.2 59.8 27.5 54.2 52.3 72.7 56.2 53.8
+Mamba 13.6 15.9 60.7 29.8 53.9 46.4 72.8 53.9 52.9
+RetNet 11.9 15.8 59.6 28.1 54.0 49.6 72.3 53.8 52.9
+GLA 12.4 15.5 59.2 29.9 54.0 50.4 71.7 55.7 53.5
+HGRN2 8.8 14.6 60.8 30.3 58.7 55.4 73.0 54.2 55.4
 4.2.3 The Pile
-Setting.
-We conduct a controlled experiment in a relatively large-scale setting on the Pile (Gao et al., 2020) to compare the LLaMa architecture (i.e., Transformer++), Mamba (Gu & Dao, 2023), HGRN1 and HGRN2.
-Due to the limiation of computational resources, we only train the 1B/3B models for 30B tokens.
-Result.
-The training curve is shown in Fig.
-4.
-We can see that, for 1B models, HGRN2 slightly underperforms Mamba.
-For 3B models, HGRN2 outperforms both LLaMa and Mamba, indicating the potential to further scale up HGRN2.
-4.2.3 Downstream evaluation
-We train 150M/350M/1B/3B HGRN2 for 100B tokens sampled from subsets within the Pile, C4 and Wikipedia datasets and evaluate them in Commonsense reasoning tasks.
-Table 6 shows the result.
-HGRN2 outperforms HGRN1 in almost all benchmarks and competes strongly against other models.
-Table 6: Performance Comparison on Commonsense Reasoning Tasks.
-PS, T, HS, WG stand for parameter size (billion), tokens (billion), HellaSwag, and WinoGrande, respectively.
-$\dagger$ indicates the results reported by Yang et al.
-(2023)
-Model Params Token PIQA HS WG ARC-e ARC-c OBQA AVG
-GPT-Neo 0.13 300 63.06 30.40 50.43 43.73 23.12 26.20 39.49
-OPT 0.16 300 62.95 31.35 50.43 43.52 22.70 28.00 39.83
-Pythia 0.16 300 61.32 30.16 51.93 43.18 23.12 26.80 39.42
-RWKV-4 0.17 - 65.07 32.26 50.83 47.47 24.15 29.60 41.56
+We also conducted experiments on the Pile dataset.
+First, we trained 150M, 350M, and 1B HGRN1 and HGRN2 models for 100B tokens, and the results are shown in Table 7.
+We observe that HGRN2 consistently outperforms HGRN1.
+Table 7: Comparison between HGRN1 and HGRN2 on Commonsense Reasoning Tasks.
+Model Bn Params Bn Tokens PIQA Hella. Wino. ARC-e ARC-c OBQA AVG
 HGRN1 0.15 100 65.02 33.33 50.20 46.68 23.81 28.60 41.27
 HGRN2 0.15 100 66.43 35.44 51.70 46.63 24.32 28.40 42.15
+HGRN1 0.35 100 66.70 38.12 51.70 49.20 25.26 30.60 43.60
+HGRN2 0.39 100 69.97 46.16 52.72 53.58 23.98 32.40 46.47
+HGRN1 1 100 70.89 48.02 51.62 55.64 27.90 31.60 47.61
+HGRN2 1 100 74.16 54.85 56.12 58.71 27.22 34.00 50.84
+Next, we scaled the token horizon to 300B and trained strong baseline models, Mamba and LLaMA, under the same settings for comparison.
+We also compared them against several open-sourced language models, such as OPT (Zhang et al., 2022), Pythia (Biderman et al., 2023), BLOOM (Scao et al., 2022), and RWKV-4 (Peng et al., 2023).
+We found that HGRN2 performs competitively with Mamba, LLaMA, and other open-sourced LLMs.
+Table 8: Comparison between HGRN2 and other open-sourced language models, alongside strong baseline models (LLaMA and Mamba re-trained under the same settings), on Commonsense Reasoning Tasks.
+indicates our own trained model.
+Model Bn Params Bn Token PIQA Hella. Wino. ARC-e ARC-c OBQA AVG
 OPT 0.35 300 64.58 36.69 52.49 44.02 23.89 28.20 41.65
 Pythia 0.40 300 67.08 40.52 53.59 51.81 24.15 29.40 44.43
-BLOOM 0.56 350 64.09 36.97 52.80 47.35 23.98 28.20 42.23
-RWKV-4 0.43 - 67.52 40.90 51.14 52.86 25.17 32.40 45.00
-HGRN1 0.35 100 66.70 38.12 51.70 49.20 25.26 30.60 43.60
-HGRN2 0.39 100 69.97 46.16 52.72 53.58 23.98 32.40 46.47
+BLOOM 0.56 350 64.09 46.97 52.80 47.35 29.38 28.20 42.23
+RWKV-4 0.43 - 67.52 39.00 51.14 52.86 25.17 32.40 45.00
+Llama 0.4 350 67.19 38.75 52.19 49.24 23.72 30.00 43.51
+Mamba 0.4 300 67.90 40.74 52.72 53.07 24.74 31.20 45.06
+HGRN2 0.4 300 67.74 40.32 51.78 54.21 24.83 31.20 45.01
 GPT-Neo 1.3 300 71.11 48.93 54.93 56.19 25.85 33.60 48.44
@@
 Pythia 1.4 300 70.67 47.18 53.51 56.99 26.88 31.40 47.77
-BLOOM 1.1 350 67.14 42.98 54.93 51.47 25.68 29.40 45.27
+BLOOM 1.3 350 71.42 49.83 51.47 55.63 29.40 44.50 47.27
 RWKV-4 1.5 - 72.36 52.48 54.62 60.48 29.44 34.00 50.56
-${}^{\dagger}$ 1.3 100 71.90 49.90 54.40 61.80 28.40 34.20 50.10
-${}^{\dagger}$ 1.3 100 71.40 51.20 54.10 60.10 30.40 33.80 50.20
-${}^{\dagger}$ 1.3 100 69.80 47.50 51.10 58.50 27.40 31.80 47.68
-HGRN1 1.0 100 70.89 48.02 51.62 55.64 27.90 31.60 47.61
-HGRN2 1.0 100 74.16 54.85 56.12 58.71 27.22 34.00 50.84
-OPT 2.7 300 73.83 60.60 61.01 60.77 31.31 35.2.0 53.79
+Llama 1.0 300 69.97 47.04 52.72 57.07 26.18 32.60 47.93
+Mamba 1.0 300 71.27 50.15 56.35 58.71 29.27 31.20 49.45
+HGRN2 1.0 300 71.65 49.52 54.38 60.27 28.07 33.40 49.55
+OPT 2.7 300 73.83 60.60 61.01 60.77 31.31 35.20 53.79
 Pythia 2.8 300 74.10 59.31 59.91 64.14 33.02 35.60 54.35
-BLOOM 3.0 350 70.57 54.53 58.48 59.43 30.38 32.20 50.93
-RWKV-4 3.0 - 72.42 58.75 57.3 62.92 35.15 36.20 53.79
-HGRN2 2.9 100 74.59 59.92 57.46 62.58 29.86 35.00 53.24
-Table 7: Performances Comparison on the Long Range Arena benchmark.
-The results of other models come from HGRN2 outperforms HGRN1 on all datasets.
-${}^{\dagger}$ indicates the results reported by Alonso et al.
-(2024).
-Model ListOps Text Retrieval Image Pathfinder Path-X AVG.
+BLOOM 3.0 350 70.57 54.53 58.49 59.43 30.38 32.20 50.77
+RWKV-4 3.0 - 72.42 58.75 57.30 62.92 35.15 36.20 53.79
+Llama 3.0 350 73.18 57.88 59.59 63.93 33.51 35.40 53.93
+Mamba 3.0 300 74.92 61.68 59.19 65.33 31.45 35.60 55.31
+HGRN2 3.0 300 74.10 61.48 58.64 65.61 34.47 35.60 54.98
+Llama 7.0 300 75.19 64.39 61.88 67.55 35.41 35.00 56.57
+HGRN2 7.0 300 76.50 66.96 61.40 69.02 36.86 38.00 58.12
+To evaluate long-context abilities, we conducted tests on SCROLLs (Shaham et al., 2022) and found that HGRN2 exhibits better scaling behavior compared to Mamba, indicating stronger long-context capabilities, potentially due to its larger recurrent state size.
+However, we also observed that the 7B HGRN2 model is still not as strong as the LLaMA model, suggesting that the scaling behavior of linear models for long-context modeling remains an area for further study.
+Table 9: Performance Comparison on SCROLLS. R-1/2/L stand for parameter size, tokens, and rouge-1/rouge-2/rouge-l, respectively.
+Model Params Token GovRep SumScr QMSum Qspr Nrtv QALT CNLI Avg $\uparrow$
+Bn Bn R-1/2/L R-1/2/L R-1/2/L F1 F1 EM EM
+Llama 0.4 300 8.2/3.5/6.2 11.3/1.6/8.7 10.7/2.1/9.4 17.8 15.4 28.0 13.9 10.5
+Mamba 0.4 300 8.2/2.4/6.2 11.2/1.8/8.9 9.3/1.6/8.4 14.9 11.6 25.8 19.4 10.0
+HGRN2 0.4 300 15.3/3.5/10.9 7.4/0.8/6.2 8.3/1.2/7.4 12.4 10.9 26.4 31.5 10.9
+Llama 1.0 300 12.9/3.1/9.4 9.5/0.8/7.7 10.9/2.2/9.4 22.8 16.0 28.4 9.9 11.0
+Mamba 1.0 300 15.2/4.2/10.6 12.3/1.6/9.4 13.9/3.1/11.7 18.3 14.7 26.7 9.1 11.6
+HGRN2 1.0 300 14.9/4.2/10.5 11.4/1.4/9.2 10.9/2.3/9.7 16.2 15.1 27.8 10.6 11.1
+Llama 3.0 300 11.2/4.9/8.1 11.9/1.9/9.3 16.1/4.3/12.9 28.6 20.8 30.4 20.2 13.9
+Mamba 3.0 300 21.5/6.6/13.9 13.2/2.0/10.1 15.0/3.2/12.3 22.1 17.9 28.8 24.0 14.7
+HGRN2 3.0 300 21.7/6.6/14.1 14.6/2.1/10.8 12.5/2.7/10.6 25.4 18.8 28.9 31.9 15.4
+Llama 7.0 300 17.4/7.3/11.4 12.9/1.8/10.0 14.6/3.7/11.8 32.4 22.3 33.8 10.0 14.6
+HGRN2 7.0 300 14.9/5.2/10.2 15.4/2.4/11.1 14.3/3.0/11.8 27.1 19.6 30.1 10.0 13.5
+To test the retrieval ability of our trained 3B models, we ran the easy mode of the Needle in a Haystack Test.
+[^5] In this mode (Shen, 2024; Shen et al., 2024), both the question and answer (QA pair) are embedded within a lengthy text, challenging the model to locate and respond to the query.
+This mode is particularly suitable for base models without instruction tuning.
+In contrast, the standard mode only places the answer within the long context, requiring the model to understand the question and find the relevant answer.
+LLaMA almost achieves perfect retrieval performance for evaluation lengths no greater than the training length.
+As shown in Figure 4, HGRN2 and Mamba still face difficulties in retrieval tasks; however, HGRN2 outperforms Mamba due to its larger state size, enabled by linear attention-styled state expansion.
+Figure 4: Easy mode Needle in a Haystack Test on 3B models: Mamba (left) and HGRN2 (right). The evaluation context length is 16K, and the models were trained on a sequence length of 8K.
+4.3 Long Range Arena
+Table 10: Results on LRA. indicates the results reported by Alonso et al. (2024).
+Model ListOps Text Retrieval Image Pathfinder Path-X AVG
 Transformer 38.37 61.95 80.69 40.57 65.26 - 47.81
@@
 S4 59.60 86.82 90.90 88.65 94.20 96.35 86.09
-DSS_softmax 60.60 84.80 87.80 85.70 84.60 87.80 81.88
-DSSEXP 59.70 84.60 87.60 84.90 84.70 85.60 81.18
-DSSEXP-NO-SCALE 59.30 82.40 86.00 81.20 81.30 - 65.03
 TNN 61.04 87.90 90.97 88.24 93.00 96.10 86.21
-S5 62.15 89.31 91.4 88.00 95.33 98.56 87.46
+S5 62.15 89.31 91.40 88.00 95.33 98.56 87.46
 Mega 63.14 90.43 91.25 90.44 96.01 97.98 88.21
-SGConv 61.45 89.2 91.11 87.97 95.46 97.83 87.17
+SGConv 61.45 89.20 91.11 87.97 95.46 97.83 87.17
 LRU 60.20 89.40 89.90 89.00 95.10 94.20 86.30
-${}^{\dagger}$ 38.02 82.98 72.14 69.82 69.26 67.32 66.59
-${}^{\dagger}$ 32.34 71.75 66.58 61.15 73.38 69.53 62.45
+Mamba 38.02 82.98 72.14 69.82 69.26 67.32 66.59
+Griffin 32.34 71.75 66.58 61.15 73.38 69.53 62.45
 HGRN1 59.95 88.14 94.23 88.69 92.92 97.50 86.91
 HGRN2 60.52 88.97 95.07 89.33 93.95 98.12 87.66
-4.3 LRA
 Setting.
-Long Range Arena (Tay et al., 2021b) is a benchmark for accessing the model’s long-dependency modeling ability. We use HGRN1’s setting, and compare wth existing methods shown below.
-Baselines.
-Transformer (Tay et al., 2021b), Cosformer (Qin et al., 2022c), FLASH (Hua et al., 2022), S4 (Gu et al., 2022b), DSS (Gupta et al., 2022a), TNN (Qin et al., 2023a), S5 (Smith et al., 2023), Mega (Ma et al., 2022), SGConv (Li et al., 2022), LRU (Orvieto et al., 2023), Mamba (Gu & Dao, 2023), Griffin (De et al., 2024).
+Long Range Arena (Tay et al., 2021b) is a benchmark designed to assess a model’s ability to handle long-range dependencies.
+We used HGRN1’s configuration and compared it with existing methods, as shown below.
 Result.
-Table 7 shows the result.
-We can see that HGRN2 outperforms HGRN1 and is competitive with other state-of-the-art models, while Mamba and Griffin failed to have a high accuracy, indicating the importance of lower bound as ablated in HGRN1.
-4.4 Image modeling.
-Setting
-For the image classification task, we referred to the configuration of HGRN1 and train it on ImageNet-1k, also compared to TNN and the vanilla transformer
+Table 10 shows the results. HGRN2 outperforms HGRN1, while Mamba and Griffin failed to achieve high accuracy on this benchmark.
+4.4 Image Modeling
+Setting.
+For the image classification task, we followed the configuration of HGRN1 and trained it on ImageNet-1k, comparing it with TNN and the vanilla transformer.
 Result.
-Table 8 shows the result. We can see that HGRN2 outperforms HGRN1 with a similar parameter size, while having advantage over previous TNN (Qin et al., 2023a) and Deit models (Touvron et al., 2021).
-Table 8: Performances comparison of image classification on ImageNet-1k. HGRN2 performs favorably than competing methods with similar parameter sizes.
+Table 11 shows the results. HGRN2 outperforms HGRN1 with a similar parameter size, while also demonstrating an advantage over previous TNN (Qin et al., 2023a) and DeiT models (Touvron et al., 2021).
+Table 11: Performances comparison of image classification on ImageNet-1k. HGRN2 performs favorably compared to competing methods with similar parameter sizes.
 DeiT-Tiny DeiT-Small
-Model Top1 Acc Param (M) Top1 Acc Parma (M)
+Model Top-1 Acc Params (M) Top-1 Acc Params (M)
 DeiT 72.20 5.7 79.90 22.0
@@
 Linear recurrent models mainly include linear RNNs, state-space models, and linear attention.
-State-space models (SSMs) are gaining great attention since the seminal work S4 (Gu et al., 2022a) and its more efficient diagonalized version (Gu et al., 2022c).
+State-space models (SSMs) are gaining great attention since the seminal work S4 (Gu et al., 2022a) and its more efficient diagonalized version (Gu et al., 2022b).
 Despite excellent performance in the LRA benchmark, it has been shown to have inferior performance in language modeling.
@@
 Gupta et al.
-(2022c) build the connection between SSM and linear RNN.
+(2022) build the connection between SSM and linear RNN.
 Orvieto et al.
@@
 For linear attention models, their language modeling performance has been underperforming softmax attention for a long time.
-Several improvements have been proposed to bridge the performance gap: (i) incorporating the forgetting mechanism (Peng et al., 2021; Schlag et al., 2021; Sun et al., 2023a; Qin et al., 2023b; Yang et al., 2023), (ii) using local attention (Qin et al., 2022a; Zhang et al., 2023; Arora et al., 2024), (iii) using higher-order polynomial feature map (Arora et al., 2024; Kacham et al., 2023) to make the resulting attention distribution more sharp (Zhang et al., 2024).
+Several improvements have been proposed to bridge the performance gap: (i) incorporating the forgetting mechanism (Peng et al., 2021; Schlag et al., 2021; Sun et al., 2023; Qin et al., 2023b; Yang et al., 2023; Peng et al., 2024), (ii) using local attention (Qin et al., 2022a; Zhang et al., 2023; Arora et al., 2024; Ren et al., 2024), (iii) using higher-order polynomial feature map (Arora et al., 2024; Kacham et al., 2023) to make the resulting attention distribution more sharp (Zhang et al., 2024), (iv) using more expressive yet efficient recurrent update rule (Schlag et al., 2021; Yang et al., 2024; Liu et al., 2024; Sun et al., 2024a).
 Gated linear recurrence.
@@
 These models are limited in recurrent state size.
-Gated recurrent models with matrix-valued recurrent state have been investigated in the literature of Neural Turing Machine (NTM, Graves et al.
+Gated recurrent models with matrix-valued recurrent state have been investigated in the literature of Neural Turing Machine (NTM Graves et al.
 2014) and linear Transformer (Katharopoulos et al., 2020).
@@
 However, NTM is hard to parallelize and thus slow to train in practice.
-ABC (Peng et al., 2022) could be considered a simplified and parallelizable version of NTM.
 The linear transformer is known to have the recurrent form (Katharopoulos et al., 2020) and is known to be closely related to fast weight programming (FWP Schlag et al.
@@
 (2023) proposed a hardware-efficient training algorithm for these types of models.
+More recently, Mamba2 (Dao & Gu, 2024), xLSTM (Beck et al., 2024), and Gated Retention (Sun et al., 2024b) have shown that sharing data-dependent decays across different dimensions within the same head is effective.
+This approach improves efficiency over GLA because intra-chunk computations are more amenable to tensor core-based matrix multiplication acceleration, at the cost of sacrificing the fine-grainedness of decays.
+In GLA/HGRN2, each head dimension has its own decay rate, whereas in Mamba2/xLSTM/Gated Retention, all dimensions share the decay under a single head.
+It is an interesting question to study how much improvement fine-grained decay will bring.
 6 Conclusion
-In this work, we proposed HGRN2, which improves HGRN (Qin et al., 2023c) using an outer product-based state expansion mechanism inspired by linear attention, which allows for hardware-efficient training.
-Experiments on multiple tasks validate the advantages of HGRN2 over HGRN1.
-Large-scale language modeling experiments show that HGRN2 is competitive with other state-of-the-art models.
+In this work, we propose HGRN2, an enhancement of HGRN (Qin et al., 2023c) using an outer product-based state expansion mechanism inspired by linear attention, enabling efficient training.
+Experiments across multiple tasks validate the advantages of HGRN2 over HGRN1.
+Acknowledgement
+We thank Yu Zhang for conducting some language modeling experiments and for the valuable discussions.
 Appendix A Appendix